Discuss how growth, risk and profitability factors cause differences in Price-to-NPAT (c) ratios across firms

You are instructed by your client to value the target company your client is considering acquiring using market-based valuation methods. Below is a summary of the financial information (dollars in millions) target company and three other companies, Company A, Company B, and Company C that you have identified as potentially comparable companies (a) as at, and for the year ended, 31 March 2018 for the Comparable companies Company Company Company C Target A Company value Equity value Net operating assets (NOA) Book value of equity Net operating profit after tax (NOPAT) Net income (NPAT) Number of shares outstanding (million) 606 2,880 1,750 1,561 1,000 ? 500 2,465 1,900 250 1,650 1,500 110 1,200 800 225 80 65 125 200 72 60 15 50 250 150 200 REQUIRED: (i) Compute the per share value of the equity for the target company as at 31 March 2018 using (1) the Price-to-NOPAT multiple, and (2) the Price-to-Book Value multiple. Round to two (2) decimal places.
(ii You found that the valuation report prepared by the target company valued the target company's equity as at 31 March 2018 at $39.50 per share. In addition, the target company's valuation report assumed the company's future ROE to remain at the current level and its discount rate to be 10%. Compute the future growth rate implied by the target company's valuation report. Round to two (2) decimal places
Given the following information (dollars in millions) for Theta Limited and Echo Limited, determine the difference in the two companies' theoretically correct Price-to- Book Value ratio at the end of 2018. Round to two (2) decimal places. Theta Echo Limited Limited Information extracted from the financial statements for the year 2018 $515 $350 Net Operating Assets (NOA) Book value of equity $515 $350 Information estimated for the terminal period starting from 2019 RNOA (based on beginning balance of NOA) 17% 13% Expected ROE 17% 13% Weighted average cost of capital 10% 10% Expected growth rate 3% 3%
Discuss how growth, risk and profitability factors cause differences in Price-to-NPAT (c) ratios across firms
Answer #1

a. Market price per share(Given)=$39.5 per share

NOPAT per share= NOPAT/No. of shares

= 225/200

=$1.13 per share

Book value of equity per share = Book value of equity/No. of shares

=1500/200

=$7.50 per share

(i)(1) Price to NOPAT multiple= Market price per share/NOPAT per share

=39.50/1.125

=35.11 times

(i)(2) Price to Book value multiple= Market price per share/ Book value per share

=39.50/7.50

=5.27 times

(ii)

ROE= Net Income/Book value of Equity*100

=200/1500*100

= 13.33%

Capitalization Rate i.e. Discount rate (given)=10%

Implied Growth rate= ROE-Capitalization Rate

=13.33-10.00

=3.33%